08:30
Beam ON ~1.0E7 pps
Target ON and open to the ring
YR09xxx ETarget ~1.9e-10 (saturated)
Target E4 inlet pressure ~8.8E-9 mbar
Dump S3 ~ 6.8E-10 mbar
Nozzle temperature 150 K
Run 54 started, monitoring code restarted.
FH and ML are optimising the beam. |